I worked this problem a million times and can get the right answer can anyone help?
(g-8)+4=-12+3g

Found 2 solutions by stanbon, chitra:
Answer by stanbon(57361)

(g-8)+4=-12+3g
g-4=-12+3g
Add 12 to both sides and subtract g from both sides to get:
8=2g
Divide both sides by 2 to get:
b=4
Cheers,
Stan H.

Answer by chitra(359)

The given expression is:

(g-8) + 4 = -12 + 3g

To solve this equation for "g", we shall first remove the brackets, and then simplify them by grouping all the like terms.

g - 8 + 4 = -12 + 3g

g - 4 = -12 + 3g

Subtracting (12 - g) from both the sides:


12 - 4 = 3g - g

8 = 2g

Therefore, g = 8/2

g = 2

Hence, the solution.

Happy calculating!!!!!!...